# Uni-X: Mitigating Modality Conflict with a Two-End-Separated Architecture for Unified Multimodal Models

### v2

@@ 1 Introduction
Vision-Language Models (VLMs) have demonstrated remarkable progress in multimodal understanding and reasoning, enabled by combining Large Language Models (LLMs) with powerful visual encoders (Liu et al., 2024c; b; Wang et al., 2024b; Team et al., 2024b). Motivated by this success, recent research has sought to extend VLMs with image generation capabilities, resulting in the development of Unified Multimodal Models (UMMs) (Team, 2025; Wang et al., 2024d; Wu et al., 2024b). However, many state-of-the-art UMMs rely on increasingly complex system designs to boost performance, including the addition of semantic image encoders (Wu et al., 2024a; Chen et al., 2025b; Deng et al., 2025; Wu et al., 2025a), the hybridization of autoregressive and diffusion paradigms (Wu et al., 2025a; Zhao et al., 2024; Ge et al., 2025; Deng et al., 2025; Xie et al., 2025; Zhou et al., 2024), or the introduction of task-specific branches and experts (Deng et al., 2025; Liao et al., 2025; Li et al., 2025c). While effective, this added complexity hinders scalability, limiting the degree of parameter sharing and reducing the potential for mutual benefits across tasks and modalities.
In contrast, autoregressive (AR) UMMs offer a simple yet powerful alternative. By treating visual inputs as a “foreign language” through vector quantization (VQ) (van den Oord et al., 2018; Esser et al., 2021b), they unify text and vision into a consistent token sequence, naturally extending the language-centric paradigm of LLMs (Wu et al., 2025b; Wang et al., 2024d). Despite this simplicity, our experiments reveal a fundamental challenge: fully modality-shared transformers trained jointly on multimodal inputs exhibits severe gradient conflicts. Originally studied in multi-task learning (Yu et al., 2020; Shi et al., 2023), we are the first to transfer this concept to UMMs, uncovering inter-modality conflicts that hinder convergence and performance.
@@
As illustrated in Figure 1, these conflicts are most pronounced in the shallow (input) and deep (output) layers, where the model must reconcile the vastly different statistical properties of text and images. In contrast, the middle layers, where representations become increasingly abstract and semantic (Meng et al., ; Geva et al., 2021; Sun et al., 2025), show reduced conflicts and stronger cross-modal alignment. This suggests that an effective UMM should respect modality-specific differences rather than enforcing uniform parameter sharing across all layers.
@@ 2 Related Work
To enable such synergy, recent efforts have shifted toward UMMs, which aim to support both understanding and generation within a single framework (Team, 2025; Jin et al., 2024; Wu et al., 2024b). A natural extension is to adopt the autoregressive (AR) paradigm of LLMs by treating visual tokens as a “foreign language” via vector quantization (Wu et al., 2025b; Wang et al., 2024c). However, the distinct statistical properties of text and images often lead to modality conflicts, degrading performance in shared transformers (Team, 2025).
To mitigate this, several approaches increase architectural complexity: Mixture-of-Transformers (MoT) designs (Liao et al., 2025; Deng et al., 2025; Shi et al., 2025) separate understanding and generation with distinct branches; Hybrid AR–diffusion frameworks (Zhao et al., 2024; Wu et al., 2025a; Dong et al., 2023; Ge et al., 2025) combine next-token prediction with diffusion-based image synthesis; and branching strategies such as UniFork (Li et al., 2025c) add task-specific deep heads. While effective on benchmarks, these methods sacrifice parameter sharing, complicate training, and weaken cross-modal benefits–the very goals UMMs were meant to unify.
@@
Uni-X builds on these insights but takes a different path. Instead of adding modules, Uni-X retains the simplicity of pure AR UMMs while mitigating modality conflict through a two-end-separated, middle-shared architecture: shallow and deep layers use modality-specific parameters to process low-level statistical differences, while intermediate layers are shared to exploit high-level semantic alignment. This X-shaped design avoids the rigidity of MoT or UniFork and the complexity of AR-diffusion hybrids, offering a lightweight yet effective solution. Empirically, Uni-X achieves comparable or superior performance to larger 7B AR UMMs across text, vision understanding, and generation tasks, while maintaining parameter efficiency and training simplicity.
@@ 3.1 Observations
However, since transformer layers have inherently different roles across depth (Sun et al., 2025; Geva et al., 2021), the resulting raw similarity $S_{\texttt{inter}}$ is biased and cannot be directly compared. To correct for this, we estimate a baseline similarity $S_{\texttt{base}}$ that reflects the model’s intrinsic gradient consistency on a unified data distribution. We randomly shuffle the full multimodal dataset $D_{\texttt{all}}$ and split it into two disjoint halves, $D_{\texttt{any}}^{1}$ and $D_{\texttt{any}}^{2}$. Their respective average gradients, $\bm{g}_{\texttt{any}}^{1}$ and $\bm{g}_{\texttt{any}}^{2}$, yield:
@@
This value represents the expected gradient similarity when gradients originate from the same underlying distribution. Therefore, we define the gradient conflict $c_{g}$ as the deviation from this baseline:
@@
To explain these observations, we examine whether vision behaves like a "foreign language" when tokenized. Using the VQ tokenizer (Team, 2025), images are represented as discrete token sequences, formally similar to text. Then, we define conditional entropy based on $n$-gram. When $n=1$, the calculation reduces to ordinary information entropy. For $n>1$, the conditional entropy is computed as follows:
@@ 3.2 Model Architecture
Importantly, unlike other architectures (Li et al., 2025c; Deng et al., 2025; Shi et al., 2025), the vision and text modalities remain strictly isolated within the separated blocks, with no cross-modal interaction. This forces the model to learn robust unimodal representations before they are fused in the shared block and after they are separated for modality-specific output generation.
@@
Unlike prior architectures that rely on auxiliary semantic encoders (Wu et al., 2024a; Chen et al., 2025b), hybrid AR–diffusion pipelines (Wu et al., 2025a; Zhao et al., 2024), or task-specific branching structures such as MoT (Liao et al., 2025) and UniFork (Li et al., 2025c), Uni-X maintains the simplicity of a pure autoregressive framework. Its two-end-separated, middle-shared structure is motivated directly by empirical evidence of gradient conflicts, aligning the model design with the statistical characteristics of each modality. By isolating low-level modality-specific processing while preserving a shared semantic core, Uni-X avoids the complexity and overhead of multi-expert or dual-paradigm systems, yet achieves competitive or superior performance. This balance of architectural simplicity, empirical grounding, and scalability makes Uni-X a practical foundation for unified multimodal modeling.
@@ 4.1 Experimental Setup
Our pre-training stage was designed to build a strong foundation in both language and vision. To preserve general text generation capabilities, we utilized a diverse set of text corpora: the high-quality Chinese dataset CCI3-H (Wang et al., 2024a), English datasets DCLM (Li et al., 2025b) and Fineweb-Edu (Penedo et al., 2024), and the StarcoderData (Li et al., 2023b) corpus, as integrating code is known to boost general model performance (MA et al., 2024). For multimodal pre-training, we used public benchmarks like ImageNet (Russakovsky et al., 2015) and JourneyDB (Pan et al., 2023), complemented by a substantial internally collected dataset of 40 million images, which were captioned using the powerful Intern-VL model Chen et al. (2024). Following the methodology of Liquid (Wu et al., 2025b), we diversified our training data by randomly reversing 20% of the text-to-image pairs to serve as image-captioning tasks. The final pre-training data consists of 72B text tokens and 65B vision tokens.
@@
We further refined the model with 3B SFT tokens. For vision understanding, we employed MiniGemini (Li et al., 2024) and FineVision (HuggingFaceM4, 2025). To improve text understanding and general instruction-following, we utilized OpenOrca (Mukherjee et al., 2023). Additionally, to refine the quality of image generation, we leveraged Blip3o-60k (Chen et al., 2025a) and ShareGPT4o (Chen et al., 2023).
@@
We conducted ablation studies on Qwen2.5-1.5B (Yang et al., 2024) and scaled to Qwen2.5-3B. We used the VQGAN tokenizer (Esser et al., 2021a) from Chameleon (Team, 2025) to encode $512\times 512$ images into $32\times 32$ discrete tokens. Our codebase is built upon the Liquid (Wu et al., 2025b) and HuggingFace Transformers (Wolf et al., 2019) libraries. Training was accelerated using Flash Attention 2 (Dao et al., 2022) and DeepSpeed ZeRO2 (Aminabadi et al., 2022). When generating images, we uniformly set the classifier-free guidance (CFG) to 4.0.
@@
Table 3: Image edit results on ImgEdit-Bench. indicates that it has been trained on more image-text data.
@@ 4.2 Results and Analysis
For image generation, as detailed in Table 2, Uni-X achieves a strong score of 82 on GenEval, a result that surpasses many models with more parameters and underscores the effectiveness of our architecture in producing high-quality images. We also tested T2I-CompBench (Huang et al., 2025) and MSCOCO (Lin et al., 2015), as shown in Appendix A.3. Uni-X similarly exhibited strong performance with fewer parameters. Regarding vision understanding (Table 2), while Uni-X’s scores are slightly lower than some state-of-the-art models, we observe a clear trend: models that incorporate an additional semantic image encoder, such as Janus-Pro (Chen et al., 2025b) and the semantically aligned variants of Liquid (Wu et al., 2025b) and Show-o (Xie et al., 2024), tend to achieve substantially higher performance on understanding benchmarks like MMBench and SEED.
In contrast, among models that do not rely on a separate semantic encoder, Uni-X’s performance is commendable and holds its ground against strong competitors like EMU3 (Wang et al., 2024d). This suggests that our architecture effectively harnesses the inherent capabilities of the autoregressive framework for vision understanding. We speculate that the relatively weaker understanding performance might be partially caused by the insufficient utilization of the VQ tokenizer’s codebook. We analyzed the token sequences encoded from 1 million images and found that, although there are 8,192 tokens available in the tokenizer, only $\approx$3,127 are being utilized. Meanwhile, EMU3 uses $4096$ tokens to represent a $512\times 512$ image, which provides more fine-grained information. However, this $4\times$ token count severely impacts its image generation speed, as shown in Appendix A.5.
@@
To validate the effectiveness of Uni-X, we conducted ablation experiments on a smaller dataset and a slightly reduced base model Qwen2.5-1.5B, due to resource constraints. To ensure consistency in performance comparisons, we limited the dataset to 2828B tokens, of which 13.713.7B are vision tokens. The experiments were conducted using learning rate (LR) $5\times 10^{-5}$, warmup ratio 0.03, and constant LR scheduler, with batch size 17,560 tokens per GPU.
The selected baselines include: (1) Shared Transformer, which continues multimodal pre-training based on Qwen2.5-1.5B; (2) Mixture-of-Transformers (MoT) (Deng et al., 2025), where prior work replicates an additional transformer to handle image generation tasks, while the original LLM backbone focuses on text-only and image understanding tasks. Under our experimental setup, vision tokens are allocated to the duplicated transformer; (3) Hard-Route MoE (HardMoE), which introduces a vision expert specifically for the vision modality, assigning vision tokens to this expert for computation guided by the vision mask; and (4) UniFork (Li et al., 2025c), which creates a task-specific deep branch for image generation. For all the baselines, we ignored the instruction during training to enhance cross-modal performance and ensure a fair comparison.
@@
In Figure  5, we present a curated selection of images generated by Uni-X to qualitatively assess its capabilities. Despite the relatively limited number of tokens used during training, the model demonstrates a strong ability to produce clear, aesthetically pleasing images that exhibit robust instruction-following capabilities. The examples showcase Uni-X’s versatility in handling a wide range of creative and complex prompts. For instance, the model can generate imaginative fantasy scenes, such as a gigantic library floating above the clouds, and surreal compositions, like a realistic elephant walking on the ocean floor. Furthermore, Uni-X successfully adheres to specific artistic style requests, as seen in the detailed anime-style portrait, and renders fine details with high fidelity, exemplified by the intricate feather patterns of the owl. These case studies collectively verify that Uni-X can effectively translate complex textual descriptions into high-quality visual outputs. The specific prompts used for these generations are provided in Appendix A.1.
@@
The results demonstrate that Uni-X can successfully interpret the contextual examples and apply the learned pattern to the target image. For instance, in the top row of Figure 6, the model correctly identifies the weather in the target image as a “rainy day,” adhering to the simple descriptive format (“This is a… day.”) established by the preceding examples. Also, Uni-X exhibits the ability to perform more reasoning tasks such as object counting. This suggests that the model is not merely mimicking sentence structure but is performing cross-modal reasoning at a semantic level.
@@
Ignoring the loss of the instruction part during training is a common technique in supervised fine-tuning. However, its role in pretraining is rarely emphasized. Following Liquid (Wu et al., 2025b), we applied the same “ignore instruction” strategy during pretraining.
Specifically, no loss mask was applied for pure text data. For text-image pairs, in text-to-image tasks, the loss calculation excluded the text instruction tokens. Similarly, for image captioning tasks, the loss corresponding to the image tokens was masked. As demonstrated in our experimental results Table 4, this approach significantly enhanced the model’s capability to generate images.
@@
We investigate how the number and distribution of separated layers affect performance (Table 5). Varying the total number of separated layers produces an $n$-shaped trend: more separation improves modality-specific low-level processing, but too many reduce shared middle layers, weakening semantic fusion and cross-modal reasoning. The best overall performance is achieved with 14 separated layers. We then examine shallow-deep ratios under this setting. A 9:5 split (slightly more shallow than deep layers) performs best, indicating that early processing of low-level features, where text and vision differ most, benefits more from modality-specific capacity than the final generation stage. These results provide strong empirical support for the Uni-X design. We also explored text layers and vision layers with different numbers of separate layers, and the results are shown in Appendix A.6.

### v3

# arXiv:2509.24365v3[cs.CV] 02 Mar 2026

@@ 1 Introduction
Vision-Language Models (VLMs) have demonstrated remarkable progress in multimodal understanding and reasoning, enabled by combining Large Language Models (LLMs) with powerful visual encoders (Liu et al., 2024c; b; Wang et al., 2024b; Team et al., 2024b). Motivated by this success, recent efforts to add image generation have led to Unified Multimodal Models (UMMs) (Team, 2025; Wang et al., 2024c; Wu et al., 2024b). However, many advanced UMMs rely on increasingly complex system designs to boost performance, including the addition of semantic image encoders (Wu et al., 2024a; Chen et al., 2025b; Deng et al., 2025; Wu et al., 2025a; Lin et al., 2025), the hybridization of autoregressive and diffusion paradigms (Wu et al., 2025a; Zhao et al., 2024; Ge et al., 2025; Deng et al., 2025; Xie et al., 2025; Zhou et al., 2024), or the introduction of task-specific branches and experts (Deng et al., 2025; Liao et al., 2025; Li et al., 2025d). While effective, this added complexity hinders scalability, limiting the degree of parameter sharing and reducing the potential for mutual benefits across tasks and modalities.
In contrast, autoregressive (AR) UMMs offer a simple yet powerful alternative. By treating visual inputs as a “foreign language” through vector quantization (VQ) (van den Oord et al., 2018; Esser et al., 2021b), they unify text and vision into a consistent token sequence, naturally extending the language-centric paradigm of LLMs (Wu et al., 2025c; Wang et al., 2024c). Despite this simplicity, our experiments reveal a fundamental challenge: fully modality-shared transformers trained jointly on multimodal inputs exhibit severe gradient conflicts. Originally studied in multi-task learning (Yu et al., 2020; Shi et al., 2023), we are the first to transfer this concept to UMMs, uncovering inter-modality conflicts that hinder convergence and performance.
@@
As illustrated in Figure 1, these conflicts are most pronounced in shallow (input) and deep (output) layers due to differing text-image statistics. In contrast, the middle layers, where representations become increasingly abstract and semantic (Meng et al., 2022; Geva et al., 2021a; Sun et al., 2025), show reduced conflicts and stronger cross-modal alignment. This suggests that an effective UMM should respect modality-specific differences rather than enforcing uniform parameter sharing across all layers.
@@ 2 Related Work
To enable such synergy, recent efforts have shifted toward UMMs, which aim to support both understanding and generation within a single framework (Team, 2025; Jin et al., 2024; Wu et al., 2024b). A natural extension is to adopt the autoregressive (AR) paradigm of LLMs by treating visual tokens as a “foreign language” via vector quantization (Wu et al., 2025c; Wang et al., 2024c). However, the distinct statistical properties of text and images often lead to modality conflicts, degrading performance in shared transformers (Team, 2025).
To mitigate this, several approaches increase architectural complexity: Mixture-of-Transformers (MoT) designs (Liao et al., 2025; Deng et al., 2025; Shi et al., 2025a) separate understanding and generation with distinct branches; Hybrid AR–diffusion frameworks (Zhao et al., 2024; Wu et al., 2025a; Dong et al., 2023; Ge et al., 2025) combine next-token prediction with diffusion-based image synthesis; and branching strategies such as UniFork (Li et al., 2025d) add task-specific deep heads. While effective on benchmarks, these methods sacrifice parameter sharing, complicate training, and weaken cross-modal benefits–the very goals UMMs were meant to unify.
@@
Uni-X builds on these insights but takes a different path. Instead of adding modules, Uni-X retains the simplicity of pure AR UMMs while mitigating modality conflict through a two-end-separated, middle-shared architecture: shallow and deep layers use modality-specific parameters to process low-level statistical differences, while intermediate layers are shared to exploit high-level semantic alignment. This X-shaped design avoids the rigidity of MoT or UniFork and the complexity of AR-diffusion hybrids, offering a lightweight yet effective solution. We further discuss other related works in Appendix A.1.
@@ 3.1 Observations
However, since transformer layers have distinct roles at depth (Sun et al., 2025; Geva et al., 2021b), the resulting raw similarity $S_{\texttt{inter}}$ is biased and cannot be directly compared. To correct for this, we estimate a baseline similarity $S_{\texttt{base}}$ that reflects the model’s intrinsic gradient consistency on a unified data distribution. We randomly shuffle the full multimodal dataset $D_{\texttt{all}}$ and split it into two disjoint halves, $D_{\texttt{any}}^{1}$ and $D_{\texttt{any}}^{2}$. Their respective average gradients, $\bm{g}_{\texttt{any}}^{1}$ and $\bm{g}_{\texttt{any}}^{2}$, yield:
@@
This value represents the expected gradient similarity when gradients originate from the same underlying distribution. Thus, we define the gradient conflict $c_{g}$ as the deviation from this baseline:
@@
To explain these observations, we examine whether vision behaves like a “foreign language” when tokenized. Using the VQ tokenizer (Team, 2025), images are represented as discrete token sequences, formally similar to text. Then, we define conditional entropy based on $n$-gram. When $n=1$, the calculation reduces to ordinary information entropy. For $n>1$, the conditional entropy is computed as follows:
@@ 3.2 Model Architecture
Importantly, unlike other architectures (Li et al., 2025d; Deng et al., 2025; Shi et al., 2025a), the vision and text modalities remain strictly isolated within the separated blocks, with no cross-modal interaction. This forces the model to learn robust unimodal representations before they are fused in the shared block and after they are separated for modality-specific output generation.
@@
Unlike prior architectures that rely on auxiliary semantic encoders (Wu et al., 2024a; Chen et al., 2025b), hybrid AR–diffusion pipelines (Wu et al., 2025a; Zhao et al., 2024), or task-specific branching structures such as MoT (Liao et al., 2025) and UniFork (Li et al., 2025d), Uni-X maintains the simplicity of a pure autoregressive framework. Its two-end-separated, middle-shared structure is motivated directly by empirical evidence of gradient conflicts, aligning the model design with the statistical characteristics of each modality. By isolating low-level modality-specific processing while preserving a shared semantic core, Uni-X avoids the complexity and overhead of multi-expert or dual-paradigm systems, yet achieves competitive or superior performance. This balance of architectural simplicity, empirical grounding, and scalability makes Uni-X a practical foundation for unified multimodal modeling.
@@ 4.1 Experimental Setup
Our pre-training stage was designed to build a strong foundation in both language and vision. To preserve general text generation capabilities, we utilized a diverse set of text corpora: the high-quality Chinese dataset CCI3-H (Wang et al., 2024a), English datasets DCLM (Li et al., 2025c) and Fineweb-Edu (Penedo et al., 2024), and the StarcoderData (Li et al., 2023b) corpus, as integrating code is known to boost general model performance (MA et al., 2024). For multimodal pre-training, we used public benchmarks like ImageNet (Russakovsky et al., 2015) and JourneyDB (Pan et al., 2023), complemented by a substantial internally collected dataset of 40 million images, which were captioned using the powerful Intern-VL model Chen et al. (2024). Following the methodology of Liquid (Wu et al., 2025c), we diversified our training data by randomly reversing 20% of the text-to-image pairs to serve as image-captioning tasks. The final pre-training data consists of 72B text tokens and 65B vision tokens.
@@
We further refined the model with 3B SFT tokens. For vision understanding, we employed MiniGemini (Li et al., 2024a) and FineVision (HuggingFaceM4, 2025). To improve text understanding and general instruction-following, we utilized OpenOrca (Mukherjee et al., 2023). Additionally, to refine the quality of image generation, we leveraged Blip3o-60k (Chen et al., 2025a) and ShareGPT4o (Chen et al., 2023).
@@
We conducted ablation studies on Qwen2.5-1.5B (Yang et al., 2024) and scaled to Qwen2.5-3B. We used the VQGAN tokenizer (Esser et al., 2021a) from Chameleon (Team, 2025) to encode $512\times 512$ images into $32\times 32$ discrete tokens. Our codebase is built upon the Liquid (Wu et al., 2025c) and HuggingFace Transformers (Wolf et al., 2019) libraries. Training was accelerated using Flash Attention 2 (Dao et al., 2022) and DeepSpeed ZeRO2 (Aminabadi et al., 2022). When generating images, we uniformly set the classifier-free guidance (CFG) to 4.0.
@@
Table 3: Image edit results on ImgEdit-Bench. denotes models trained on extra image-text data.
@@ 4.2 Results and Analysis
For image generation, as detailed in Table 2, Uni-X achieves a strong score of 82 on GenEval, a result that surpasses many models with more parameters and underscores the effectiveness of our architecture in producing high-quality images. We also tested T2I-CompBench (Huang et al., 2025) and MSCOCO (Lin et al., 2015), as shown in Appendix A.4. Uni-X similarly exhibited strong performance with fewer parameters. Regarding vision understanding (Table 2), while Uni-X’s scores are slightly lower than some state-of-the-art models, we observe a clear trend: models that incorporate an additional semantic image encoder, such as Janus-Pro (Chen et al., 2025b) and the semantically aligned variants of Liquid (Wu et al., 2025c) and Show-o (Xie et al., 2024), tend to achieve substantially higher performance on understanding benchmarks like MMBench and SEED.
In contrast, among models that do not rely on a separate semantic encoder, Uni-X’s performance is commendable and holds its ground against strong competitors like EMU3 (Wang et al., 2024c). This suggests that our architecture effectively harnesses the inherent capabilities of the autoregressive framework for vision understanding. We speculate that the relatively weaker understanding performance might be partially caused by the insufficient utilization of the VQ tokenizer’s codebook. We analyzed the token sequences encoded from 1 million images and found that, although there are 8,192 tokens available in the tokenizer, only $\approx$3,127 are being utilized. Meanwhile, EMU3 uses $4096$ tokens to represent a $512\times 512$ image, which provides more fine-grained information. However, this $4\times$ token count severely impacts its image generation speed, as shown in Appendix A.6.
@@
To validate the effectiveness of Uni-X, we conducted ablation experiments on a smaller dataset and a slightly reduced base model Qwen2.5-1.5B, due to resource constraints. To ensure consistency in performance comparisons, we limited the dataset to 2828B tokens, of which 13.713.7B are vision tokens. The experiments were conducted using a learning rate (LR) of $5\times 10^{-5}$, warmup ratio 0.03, and constant LR scheduler, with batch size 17,560 tokens per GPU.
The selected baselines include: (1) Shared Transformer, which continues multimodal pre-training based on Qwen2.5-1.5B; (2) Mixture-of-Transformers (MoT) (Deng et al., 2025), where prior work replicates an additional transformer to handle image generation tasks, while the original LLM backbone focuses on text-only and image understanding tasks. Under our experimental setup, vision tokens are allocated to the duplicated transformer; (3) Hard-Route MoE (HardMoE), which introduces a vision expert specifically for the vision modality, assigning vision tokens to this expert for computation guided by the vision mask; and (4) UniFork (Li et al., 2025d), which creates a task-specific deep branch for image generation. For all the baselines, we ignored the instruction during training to enhance cross-modal performance and ensure a fair comparison.
@@
In Figure  5, we present a curated selection of images generated by Uni-X to qualitatively assess its capabilities. Despite the relatively limited number of tokens used during training, the model demonstrates a strong ability to produce clear, aesthetically pleasing images that exhibit robust instruction-following capabilities. The examples showcase Uni-X’s versatility in handling a wide range of creative and complex prompts. For instance, the model can generate imaginative fantasy scenes, such as a gigantic library floating above the clouds, and surreal compositions, like a realistic elephant walking on the ocean floor. Furthermore, Uni-X successfully adheres to specific artistic style requests, as seen in the detailed anime-style portrait, and renders fine details with high fidelity, exemplified by the intricate feather patterns of the owl. These case studies verify that Uni-X can effectively translate complex textual descriptions into high-quality visual outputs. The specific prompts used for these generations are provided in Appendix A.2.
@@
The results demonstrate that Uni-X can successfully interpret the contextual examples and apply the learned pattern to the target image. For instance, in the top row of Figure 6, the model correctly identifies the weather in the target image as a “rainy day,” adhering to the simple descriptive format (“This is a… day.”) established by the preceding examples. Also, Uni-X exhibits the ability to perform more reasoning tasks, such as object counting. This suggests that the model is not merely mimicking sentence structure but is performing cross-modal reasoning at a semantic level.
@@
Ignoring the loss of the instruction part during training is a common technique in supervised fine-tuning. However, its role in pretraining is rarely emphasized. Following Liquid (Wu et al., 2025c), we applied the same “ignore instruction” strategy during pretraining.
Specifically, no loss mask was applied for pure text data. For text-image pairs, in text-to-image tasks, the loss calculation excluded the text instruction tokens. Similarly, for image captioning tasks, the loss corresponding to the image tokens was masked. As demonstrated in our experimental results, Table 4, this approach significantly enhanced the model’s capability to generate images.
@@
We investigate how the number and distribution of separated layers affect performance (Table 5). Varying the total number of separated layers produces an $n$-shaped trend: more separation improves modality-specific low-level processing, but too many layers reduce shared middle layers, weakening semantic fusion and cross-modal reasoning. The best overall performance is achieved with 14 separated layers.
We then examine shallow-deep ratios under this setting. A 9:5 split (slightly more shallow than deep layers) performs best, indicating that early processing of low-level features, where text and vision differ most, benefits more from modality-specific capacity than the final generation stage. These results provide strong empirical support for the Uni-X design. We also explored text layers and vision layers with different numbers of separate layers, and the results are shown in Appendix A.7.